# QWHA: Quantization-Aware Walsh-Hadamard Adaptation for Parameter-Efficient Fine-Tuning on Large Language Models

### v3

@@ Abstract
The demand for efficient deployment of large language models (LLMs) has driven interest in quantization, which reduces inference cost, and parameter-efficient fine-tuning (PEFT), which lowers training overhead. This motivated the development of quantization-aware PEFT to produce accurate yet efficient quantized models. In this setting, reducing quantization error prior to fine-tuning is crucial for achieving high model accuracy. However, existing methods that rely on low-rank adaptation suffer from limited representational capacity. Recent Fourier-related transform (FT)-based adapters offer greater representational power than low-rank adapters, but their direct integration into quantized models often results in ineffective error reduction and increased computational overhead. To overcome these limitations, we propose QWHA, a method that integrates FT-based adapters into quantized models by employing the Walsh-Hadamard Transform (WHT) as the transform kernel, together with a novel adapter initialization scheme incorporating adaptive parameter selection and value refinement. We demonstrate that QWHA effectively mitigates quantization errors while facilitating fine-tuning, and that its design substantially reduces computational cost. Experimental results show that QWHA consistently outperforms baselines in low-bit quantization accuracy and achieves significant training speedups over existing FT-based adapters. The code is available at https://github.com/vantaa89/qwha.
@@ 1 Introduction
Fine-tuning enables large language models (LLMs) to generalize beyond their pre-training, allowing adaptation to various domains (Wei et al., 2022; Liu et al., 2023; Qin et al., 2024; DeepSeek-AI et al., 2025). While full fine-tuning yields superior accuracy, it often incurs significant overhead due to the extensive computations required to update all the trainable model parameters (Loshchilov & Hutter, 2017; Zhu et al., 2025). Parameter-efficient fine-tuning (PEFT) addresses this issue by optimizing only a small subset of the parameters while leaving most of them frozen (Li & Liang, 2021; Liu et al., 2022; Hu et al., 2022; Liu et al., 2024; Kopiczko et al., 2024). Beyond reducing training overhead, recent studies have shown that combining PEFT with model compression techniques can enhance inference efficiency at the same time (Dettmers et al., 2023). Among these techniques, quantization, which lowers the bit precision of model parameters, has gained particular attention due to its robustness against accuracy degradation under high compression ratios (Frantar et al., 2023; Lin et al., 2024; Dettmers et al., 2024; Kim et al., 2024b; Shao et al., 2024; Ashkboos et al., 2024; Zhang et al., 2024; Liu et al., 2025). Consequently, quantization-aware PEFT (QA-PEFT) has been widely explored as a promising approach for efficient adaptation and inference in LLMs.
@@
Prior works on QA-PEFT typically relied on low-rank adaptation (LoRA) (Li et al., 2024; Guo et al., 2024; Kim et al., 2024a; Liao et al., 2024; Deng et al., 2025). In contrast, for standard PEFT, several alternatives to LoRA have recently been proposed to address the representational limitations of low-rank structures. In particular, Fourier-related transform (FT)-based adapters have emerged as strong alternatives. They train a sparse set of coefficients to represent weight updates in the transform domain, offering superior representational capacity (Gao et al., 2024b; Du et al., 2025; Shen et al., 2025). However, our observations show that directly applying FT-based adapters to quantized models often yields worse performance than LoRA-based methods specifically designed for QA-PEFT. This highlights the importance of explicit consideration for quantization effects when fine-tuning quantized models. LoRA-based methods adopt quantization-aware initialization strategies that compensate for the errors between full- and low-precision weights using low-rank approximation with the adapters prior to fine-tuning. However, applying such initialization in FT-based adapters is non-trivial, as identifying the optimal sparse set of parameters and their values to approximate a given matrix is an NP-hard problem (Natarajan, 1995). Moreover, the choice of transform type becomes an additional design consideration. This raises a research question: how to effectively exploit FT-based adapters in QA-PEFT. To the best of our knowledge, neither FT-based adapters nor their initialization techniques have been explored in the context of QA-PEFT.
@@
In this paper, we present QWHA, a novel QA-PEFT method that introduces a FT-based adapter together with a quantization-aware initialization scheme, as illustrated in Figure 1 and Table 1. We adopt WHT in our adapter design (WHA), inspired by its high-fidelity reconstruction ability in the spectral domain, to effectively compensate for quantization errors (Hedayat, 1978). In addition, the WHT kernel consists solely of $\pm 1$ elements, enabling efficient computations using only additions and subtractions, thereby eliminating matrix multiplications (Dao-AILab, 2024). We further reduce computation by applying a single transform in the adapter, unlike conventional FT-based adapters that apply two transforms. For quantization-aware adapter initialization, we develop a tractable solution that first selects parameter locations $\bm{E}$} and then assigns their values $\bm{c}$}. We introduce a channel-wise parameter allocation scheme that guarantees a lower bound on the number of parameters per channel to facilitate fine-tuning while allocating more parameters to channels with larger quantization errors, and then select the highest-magnitude coefficients within each channel to effectively reduce quantization error (AdaAlloc). Finally, we refine the selected parameter values, thereby enabling substantial reduction of quantization error (Refinement). We theoretically analyze the superior representation capacity of our proposed adapter and empirically validate the benefits of our adapter design and initialization method across diverse datasets and models.
@@ 2.1 LLM Quantization
LLM quantization is a key technique for improving inference efficiency by reducing the memory bottleneck caused by model weights through lowering their bit precision (Frantar et al., 2023), typically expressed by the following equation:
@@
LLMs generally contain outliers, a small fraction of weights that are exceptionally large compared to the main distribution, and LLM quantization is highly sensitive to these outliers (Dettmers et al., 2024; Kim et al., 2024b; Tseng et al., 2024; An et al., 2025). These outliers induce corresponding outliers in the quantization error. Most quantization errors $\Delta{\bm{W}}_{Q}={\bm{W}}_{0}-{\bm{W}}_{Q}$ are bounded within a small range (e.g., $\tfrac{s}{2},\tfrac{s}{2})$), since most weights within the clamping range are mapped to the nearest quantization level. In contrast, for outliers, the quantization error is defined as the difference between the original large weight and the clamping boundary values, resulting in extremely large errors that lead to significant accuracy degradation. Thus, reducing outlier-induced error is critical, and recent post-training quantization techniques for LLMs focus on mitigating these errors to preserve model accuracy (Dettmers et al., 2024; Kim et al., 2024b; Shao et al., 2024; Tseng et al., 2024; Zhang et al., 2024). Details on the distribution of quantization errors are presented in Appendix A.
@@ 2.2 Quantization-Aware PEFT
A typical quantization-aware PEFT (QA-PEFT) adopts LoRA (Hu et al., 2022), which injects a pair of low-rank matrices into linear layers to approximate the weight updates $\Delta{\bm{W}}$ as follows:
@@
Here, ${\bm{A}}\in\mathbb{R}^{{r}\times d_{\text{in}}}$ and ${\bm{B}}\in\mathbb{R}^{d_{\text{out}}\times r}$ are low-rank adapters, fine-tuned instead of frozen quantized weight ${\bm{W}}_{Q}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, where ${\bm{X}}\in\mathbb{R}^{d_{\text{in}}\times(b\times s)}$ is the activation matrix with batch size $b$ and sequence length $s$. Since there is no prior information about the weight updates before fine-tuning, LoRA typically initializes $\bm{A}$} as a random matrix and $\bm{B}$} as a zero matrix. In QA-PEFT, however, initializing the adapters to minimize quantization error prior to fine-tuning plays a crucial role in accuracy. Early approaches addressed this by reconstructing quantization errors via singular value decomposition (SVD) to initialize low-rank adapters (Li et al., 2024; Guo et al., 2024). More recent works, such as RA-LoRA (Kim et al., 2024a) and CLoQ (Deng et al., 2025), adopt advanced decomposition strategies and improved calibration to further mitigate this limitation. However, existing QA-PEFT methods remain restricted to LoRA, and no prior studies have explored the use of other advanced adapters for QA-PEFT, which will be discussed in the next section.
@@ 2.3 Fourier Transform-based Adapters
Sparse adapters have recently emerged as a strong alternative to various low-rank adapters (Bhardwaj et al., 2024; Gao et al., 2024b; Shen et al., 2025; Du et al., 2025). SHiRA (Bhardwaj et al., 2024) proposes directly updating a sparse subset of the weight matrix, enabling multi-adapter fusion. More recent methods adopt Fourier-related Transforms (FT) to represent the weight update $\Delta{\bm{W}}$ in the spectral domain by applying transforms along both the rows and columns of the matrix as follows:
@@
Here, ${\bm{H}}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{in}}}$ and ${\bm{H}}^{\prime}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{out}}}$ are the orthonormal transform kernels. Prior works on these FT-based adapters have primarily focused on identifying suitable transform kernels. FourierFT (Gao et al., 2024b) employs the discrete Fourier transform (DFT), while LoCA (Du et al., 2025) replaces the DFT with the discrete cosine transform (DCT) to avoid discarding imaginary components. SSH (Shen et al., 2025) instead leverages the discrete Hartley transform (DHT) for the same purpose. As these kernels are composed of sinusoidal functions, $\bm{F}$} corresponds to the coefficients of the frequency components, which collectively represent $\Delta{\bm{W}}$. We denote DCT and DHT-based adapters as DCA and DHA throughout the paper.
Since the transform kernels are fixed matrices, $\bm{F}$} is the only learnable parameter during fine-tuning. To reduce the number of trainable parameters, $\bm{F}$} is treated as a sparse matrix. Specifically, ${\bm{F}}=\mathrm{Scatter}({\bm{c}},{\bm{E}})$ is constructed from a value vector ${\bm{c}}\in\mathbb{R}^{p}$ and an index list ${\bm{E}}\in\mathbb{N}^{p\times 2}$, where $\mathrm{Scatter}$ assigns ${\bm{F}}_{(E_{l,1},E_{l,2})}=c_{l}$ for $0\leq l\leq p-1$, with all other entries fixed to zero throughout training and inference. At the initialization stage, since there is no information on $\Delta{\bm{W}}$, previous works generally select the locations $\bm{E}$} randomly and the values of the spectral coefficients $\bm{c}$} are initialized to zero (Gao et al., 2024b; Du et al., 2025). SSH (Shen et al., 2025) proposes an advanced parameter selection strategy under the assumption that the frequency patterns of pre-trained and fine-tuned weights are similar. It first transforms the pre-trained weights and selects half of the positions with the largest spectral coefficients, while the remaining half are chosen randomly.
@@ 3.1 QA-PEFT Adapter Design
We design our proposed adapter by constructing the weight update as the transformation of a sparse matrix $\bm{F}$} through an orthogonal transform ${\bm{H}}^{-1}$. Specifically, we adopt the WHT (Hedayat, 1978; Kunz, 1979), a particular instance of the FT whose kernel consists only of $\pm 1$ entries, for the transform $\bm{H}$} (details on WHT and other FT kernels are provided in Appendix B.1). Accordingly, our adapter is formulated as follows:
@@
FT-based adapters exhibit greater representational capability than LoRA variants because they offer higher rank capacity given the same number of parameters. The representational power of low-rank adapters is strictly bounded by their inner dimension $r$ (Equation 2). In contrast, since the transform kernels in FT-based adapters are orthogonal and therefore full-rank, the rank of the adapter depends solely on the sparse matrix $\bm{F}$} (Equation 3 and 4). Given that nonzero parameters are selected uniformly at random, if both rows and columns receive more than two parameters on average, then $\bm{F}$} achieves full rank $r_{\text{max}}=\min(d_{\text{in}},d_{\text{out}})$ with high probability (Coja-Oghlan et al., 2020). Since our adapter initialization in Section 3.2 assigns at least a few elements to each channel and selects parameters independently per channel, the full-rank conditions are satisfied. Details of this condition are provided in Appendix B.2. Figure 2(a) presents the empirical analysis of the rank of adapter weights, normalized by the maximum achievable rank $r_{\text{max}}$ and averaged across layers. While LoRA achieves less than 6.3% of the normalized rank, FT-based adapters are nearly full-rank. Hence, our proposed WHA exhibits high representational capacity.
@@
As discussed in Section 2.1, quantization errors exhibit heavy-tailed outliers. For QA-PEFT, where mitigating such errors is crucial, the adapter must capture the outlier structure with a small number of parameters, as in the case of sparse adapters using the sparse matrix $\bm{F}$}. We strategically adopt the WHT for our adapter design to effectively capture such outliers (Hedayat, 1978; Kunz, 1979). The WHT kernel consists only of $\pm 1$ entries, and its basis functions are square-wave patterns with sharp transitions. In contrast, prior FT-based adapters adopt DCT or DHT, whose sinusoidal bases exhibit smooth transitions. This structural difference makes the WHT better aligned with abrupt changes such as outlier values. Therefore, WHT inherently provides a more compact coefficient representation of quantization errors compared to DCT or DHT. We empirically demonstrate this by analyzing the cumulative energy in adapter parameters (Figure 2(b)), defined as the $\ell_{2}$ norm of coefficients from the transform of $\Delta{\bm{W}}_{Q}$ in FT-based adapters, and the $\ell_{2}$ norm of singular values of $\Delta{\bm{W}}_{Q}$ in low-rank adapters. Both coefficients and singular values follow a Pareto-like distribution (see Appendix B.3), which can be characterized by the Pareto hill index $\eta$, where a smaller $\eta$ indicates a sharper distribution (Arnold, 1983). Since the total cumulative energy equals $\lVert\Delta{\bm{W}}_{Q}\rVert_{F}^{2}$, the fastest convergence curve of WHT, with the smallest $\eta$, demonstrates that it concentrates the largest portion of energy within a small number of coefficients, enabling accurate reconstruction with a limited number of parameters. As a result, WHA effectively compensates for quantization errors, particularly large-magnitude ones from salient weight channels, as shown empirically in Figure 3. For a fair comparison, we use the same parameter initialization method described in Section 3.2. We define outlier coverage as the ratio of the $\ell_{1}$ sum of coefficients captured by the selected parameter locations to that of all coefficients corresponding to the top 10% magnitude outliers of $\Delta{\bm{W}}_{Q}$.

### v4

@@ Abstract
The demand for efficient deployment of large language models (LLMs) has driven interest in quantization, which reduces inference cost, and parameter-efficient fine-tuning (PEFT), which lowers training overhead. This motivated the development of quantization-aware PEFT to produce accurate yet efficient quantized models. In this setting, reducing quantization error prior to fine-tuning is crucial for achieving high model accuracy. However, existing methods that rely on low-rank adaptation suffer from limited representational capacity. Recent Fourier-related transform (FT)-based adapters offer greater representational power than low-rank adapters, but their direct integration into quantized models often results in ineffective error reduction and increased computational overhead. To overcome these limitations, we propose QWHA, a method that integrates FT-based adapters into quantized models by employing the Walsh-Hadamard Transform (WHT) as the transform kernel, together with a novel adapter initialization scheme incorporating adaptive parameter selection and value refinement. We demonstrate that QWHA effectively mitigates quantization errors while facilitating fine-tuning, and that its design substantially reduces computational cost. Experimental results show that QWHA consistently outperforms baselines in low-bit quantization accuracy and achieves significant training speedups over existing FT-based adapters. The code is publicly available at https://github.com/vantaa89/qwha.
@@ 1 Introduction
Fine-tuning enables large language models (LLMs) to generalize beyond their pre-training, allowing adaptation to various domains (wei2022finetunedlanguagemodelszeroshot; liu2023visualinstructiontuning; qin2023toolllmfacilitatinglargelanguage; deepseekai2025deepseekr1incentivizingreasoningcapability). While full fine-tuning yields superior accuracy, it often incurs significant overhead due to the extensive computations required to update all the trainable model parameters (loshchilov2017adamw; adazhu2025apollosgdlikememoryadamwlevel). Parameter-efficient fine-tuning (PEFT) addresses this issue by optimizing only a small subset of the parameters while leaving most of them frozen (li2021prefix; liu2022fewshot; hu2022lora; liu2024dora; kopiczko2024vera). Beyond reducing training overhead, recent studies have shown that combining PEFT with model compression techniques can enhance inference efficiency at the same time (dettmers2023qlora). Among these techniques, quantization, which lowers the bit precision of model parameters, has gained particular attention due to its robustness against accuracy degradation under high compression ratios (frantar2023gptq; lin2024awqactivationawareweightquantization; dettmers2024spqr; kim2024squeezellm; shao2024omniquant; ashkboos2024quarot; zhang2024magrweightmagnitudereduction; liu2025spinquantllmquantizationlearned). Consequently, quantization-aware PEFT (QA-PEFT) has been widely explored as a promising approach for efficient adaptation and inference in LLMs.
@@
Prior works on QA-PEFT typically relied on low-rank adaptation (LoRA) (li2024loftq; guo2024lqloralowrankplusquantized; kim2024ra; liao2024apiq; deng2025cloq). In contrast, for standard PEFT, several alternatives to LoRA have recently been proposed to address the representational limitations of low-rank structures. In particular, Fourier-related transform (FT)-based adapters have emerged as strong alternatives. They train a sparse set of coefficients to represent weight updates in the transform domain, offering superior representational capacity (gao2024fourierft; du2025loca; shen2025ssh). However, our observations show that directly applying FT-based adapters to quantized models often yields worse performance than LoRA-based methods specifically designed for QA-PEFT. This highlights the importance of explicit consideration for quantization effects when fine-tuning quantized models. LoRA-based methods adopt quantization-aware initialization strategies that compensate for the errors between full- and low-precision weights using low-rank approximation with the adapters prior to fine-tuning. However, applying such initialization in FT-based adapters is non-trivial, as identifying the optimal sparse set of parameters and their values to approximate a given matrix is an NP-hard problem (natarajan1995). Moreover, the choice of transform type becomes an additional design consideration. This raises a research question: how to effectively exploit FT-based adapters in QA-PEFT. To the best of our knowledge, neither FT-based adapters nor their initialization techniques have been explored in the context of QA-PEFT.
@@
In this paper, we present QWHA, a novel QA-PEFT method that introduces a FT-based adapter together with a quantization-aware initialization scheme, as illustrated in Figure 1 and Table 1. We adopt WHT in our adapter design (WHA), inspired by its high-fidelity reconstruction ability in the spectral domain, to effectively compensate for quantization errors (hedayat1978wht). In addition, the WHT kernel consists solely of $\pm 1$ elements, enabling efficient computations using only additions and subtractions, thereby eliminating matrix multiplications (fasthadamardtransform_github). We further reduce computation by applying a single transform in the adapter, unlike conventional FT-based adapters that apply two transforms. For quantization-aware adapter initialization, we develop a tractable solution that first selects parameter locations $\bm{E}$} and then assigns their values $\bm{c}$}. We introduce a channel-wise parameter allocation scheme that guarantees a lower bound on the number of parameters per channel to facilitate fine-tuning while allocating more parameters to channels with larger quantization errors, and then select the highest-magnitude coefficients within each channel to effectively reduce quantization error (AdaAlloc). Finally, we refine the selected parameter values, thereby enabling substantial reduction of quantization error (Refinement). We theoretically analyze the superior representation capacity of our proposed adapter and empirically validate the benefits of our adapter design and initialization method across diverse datasets and models.
@@ 2.1 LLM Quantization
LLM quantization is a key technique for improving inference efficiency by reducing the memory bottleneck caused by model weights through lowering their bit precision (frantar2023gptq), typically expressed by the following equation:
@@
LLMs generally contain outliers, a small fraction of weights that are exceptionally large compared to the main distribution, and LLM quantization is highly sensitive to these outliers (dettmers2024spqr; kim2024squeezellm; tseng2024quipbetterllmquantization; an2025systematicoutlierslargelanguage). These outliers induce corresponding outliers in the quantization error. Most quantization errors $\Delta{\bm{W}}_{Q}={\bm{W}}_{0}-{\bm{W}}_{Q}$ are bounded within a small range (e.g., $\tfrac{s}{2},\tfrac{s}{2})$), since most weights within the clamping range are mapped to the nearest quantization level. In contrast, for outliers, the quantization error is defined as the difference between the original large weight and the clamping boundary values, resulting in extremely large errors that lead to significant accuracy degradation. Thus, reducing outlier-induced error is critical, and recent post-training quantization techniques for LLMs focus on mitigating these errors to preserve model accuracy (dettmers2024spqr; kim2024squeezellm; shao2024omniquant; tseng2024quipbetterllmquantization; zhang2024magrweightmagnitudereduction). Details on the distribution of quantization errors are presented in Appendix LABEL:appendix:outlier.
@@ 2.2 Quantization-Aware PEFT
A typical quantization-aware PEFT (QA-PEFT) adopts LoRA (hu2022lora), which injects a pair of low-rank matrices into linear layers to approximate the weight updates $\Delta{\bm{W}}$ as follows:
@@
Here, ${\bm{A}}\in\mathbb{R}^{{r}\times d_{\text{in}}}$ and ${\bm{B}}\in\mathbb{R}^{d_{\text{out}}\times r}$ are low-rank adapters, fine-tuned instead of frozen quantized weight ${\bm{W}}_{Q}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$, where ${\bm{X}}\in\mathbb{R}^{d_{\text{in}}\times(b\times s)}$ is the activation matrix with batch size $b$ and sequence length $s$. Since there is no prior information about the weight updates before fine-tuning, LoRA typically initializes $\bm{A}$} as a random matrix and $\bm{B}$} as a zero matrix. In QA-PEFT, however, initializing the adapters to minimize quantization error prior to fine-tuning plays a crucial role in accuracy. Early approaches addressed this by reconstructing quantization errors via singular value decomposition (SVD) to initialize low-rank adapters (li2024loftq; guo2024lqloralowrankplusquantized). More recent works, such as RA-LoRA (kim2024ra) and CLoQ (deng2025cloq), adopt advanced decomposition strategies and improved calibration to further mitigate this limitation. However, existing QA-PEFT methods remain restricted to LoRA, and no prior studies have explored the use of other advanced adapters for QA-PEFT, which will be discussed in the next section.
@@ 2.3 Fourier Transform-based Adapters
Sparse adapters have recently emerged as a strong alternative to various low-rank adapters (bhardwaj2024shira; gao2024fourierft; shen2025ssh; du2025loca). SHiRA (bhardwaj2024shira) proposes directly updating a sparse subset of the weight matrix, enabling multi-adapter fusion. More recent methods adopt Fourier-related Transforms (FT) to represent the weight update $\Delta{\bm{W}}$ in the spectral domain by applying transforms along both the rows and columns of the matrix as follows:
@@
Here, ${\bm{H}}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{in}}}$ and ${\bm{H}}^{\prime}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{out}}}$ are the orthonormal transform kernels. Prior works on these FT-based adapters have primarily focused on identifying suitable transform kernels. FourierFT (gao2024fourierft) employs the discrete Fourier transform (DFT), while LoCA (du2025loca) replaces the DFT with the discrete cosine transform (DCT) to avoid discarding imaginary components. SSH (shen2025ssh) instead leverages the discrete Hartley transform (DHT) for the same purpose. As these kernels are composed of sinusoidal functions, $\bm{F}$} corresponds to the coefficients of the frequency components, which collectively represent $\Delta{\bm{W}}$. We denote DCT and DHT-based adapters as DCA and DHA throughout the paper.
Since the transform kernels are fixed matrices, $\bm{F}$} is the only learnable parameter during fine-tuning. To reduce the number of trainable parameters, $\bm{F}$} is treated as a sparse matrix. Specifically, ${\bm{F}}=\mathrm{Scatter}({\bm{c}},{\bm{E}})$ is constructed from a value vector ${\bm{c}}\in\mathbb{R}^{p}$ and an index list ${\bm{E}}\in\mathbb{N}^{p\times 2}$, where $\mathrm{Scatter}$ assigns ${\bm{F}}_{(E_{l,1},E_{l,2})}=c_{l}$ for $0\leq l\leq p-1$, with all other entries fixed to zero throughout training and inference. At the initialization stage, since there is no information on $\Delta{\bm{W}}$, previous works generally select the locations $\bm{E}$} randomly and the values of the spectral coefficients $\bm{c}$} are initialized to zero (gao2024fourierft; du2025loca). SSH (shen2025ssh) proposes an advanced parameter selection strategy under the assumption that the frequency patterns of pre-trained and fine-tuned weights are similar. It first transforms the pre-trained weights and selects half of the positions with the largest spectral coefficients, while the remaining half are chosen randomly.
@@ 3.1 QA-PEFT Adapter Design
We design our proposed adapter by constructing the weight update as the transformation of a sparse matrix $\bm{F}$} through an orthogonal transform ${\bm{H}}^{-1}$. Specifically, we adopt the WHT (hedayat1978wht; kunz1979hadamard), a particular instance of the FT whose kernel consists only of $\pm 1$ entries, for the transform $\bm{H}$} (details on WHT and other FT kernels are provided in Appendix LABEL:appendix:wht_form). Accordingly, our adapter is formulated as follows:
@@
FT-based adapters exhibit greater representational capability than LoRA variants because they offer higher rank capacity given the same number of parameters. The representational power of low-rank adapters is strictly bounded by their inner dimension $r$ (Equation 2). In contrast, since the transform kernels in FT-based adapters are orthogonal and therefore full-rank, the rank of the adapter depends solely on the sparse matrix $\bm{F}$} (Equation 3 and 4). Given that nonzero parameters are selected uniformly at random, if both rows and columns receive more than two parameters on average, then $\bm{F}$} achieves full rank $r_{\text{max}}=\min(d_{\text{in}},d_{\text{out}})$ with high probability (amin2023rank). Since our adapter initialization in Section 3.2 assigns at least a few elements to each channel and selects parameters independently per channel, the full-rank conditions are satisfied. Details of this condition are provided in Appendix LABEL:appendix:qhft_rank. Figure 2(a) presents the empirical analysis of the rank of adapter weights, normalized by the maximum achievable rank $r_{\text{max}}$ and averaged across layers. While LoRA achieves less than 6.3% of the normalized rank, FT-based adapters are nearly full-rank. Hence, our proposed WHA exhibits high representational capacity.
@@
As discussed in Section 2.1, quantization errors exhibit heavy-tailed outliers. For QA-PEFT, where mitigating such errors is crucial, the adapter must capture the outlier structure with a small number of parameters, as in the case of sparse adapters using the sparse matrix $\bm{F}$}. We strategically adopt the WHT for our adapter design to effectively capture such outliers (hedayat1978wht; kunz1979hadamard). The WHT kernel consists only of $\pm 1$ entries, and its basis functions are square-wave patterns with sharp transitions. In contrast, prior FT-based adapters adopt DCT or DHT, whose sinusoidal bases exhibit smooth transitions. This structural difference makes the WHT better aligned with abrupt changes such as outlier values. Therefore, WHT inherently provides a more compact coefficient representation of quantization errors compared to DCT or DHT. We empirically demonstrate this by analyzing the cumulative energy in adapter parameters (Figure 2(b)), defined as the $\ell_{2}$ norm of coefficients from the transform of $\Delta{\bm{W}}_{Q}$ in FT-based adapters, and the $\ell_{2}$ norm of singular values of $\Delta{\bm{W}}_{Q}$ in low-rank adapters. Both coefficients and singular values follow a Pareto-like distribution (see Appendix LABEL:appendix:wht_energy), which can be characterized by the Pareto hill index $\eta$, where a smaller $\eta$ indicates a sharper distribution (arnold1983pareto). Since the total cumulative energy equals $\lVert\Delta{\bm{W}}_{Q}\rVert_{F}^{2}$, the fastest convergence curve of WHT, with the smallest $\eta$, demonstrates that it concentrates the largest portion of energy within a small number of coefficients, enabling accurate reconstruction with a limited number of parameters. As a result, WHA effectively compensates for quantization errors, particularly large-magnitude ones from salient weight channels, as shown empirically in Figure 3. For a fair comparison, we use the same parameter initialization method described in Section 3.2. We define outlier coverage as the ratio of the $\ell_{1}$ sum of coefficients captured by the selected parameter locations to that of all coefficients corresponding to the top 10% magnitude outliers of $\Delta{\bm{W}}_{Q}$.